OpenStudy (anonymous):

If a 0.311 M solution of a weak acid is 0.0847% ionized, what is the ionization constant for the acid?

OpenStudy (aaronq):

\(HA\rightleftharpoons H^++A^-\); \(K_A=\dfrac{[H^+][A^-]}{[HA]}\) \([H^+]=[A^-]=0.311*(0.000847)M\) \([HA]=0.311~M-0.311*(0.000847)M\) \(K_A=\dfrac{(0.311*(0.000847)M)^2}{(0.311~M-0.311*(0.000847)M)}\)